40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 177     Entry time: Mon Jun 28 19:06:40 2010
Author: tarac 
Type: DailyProgress 
Category: PMC 
Subject: PMC acting weird 

This morning, the PMC couldn't be locked to the laser. The FSS servo was disabled during that time.

When the gain/ RF power were adjusted, PMC was locked to the laser, but the coupling efficiency dropped from 80% to 40%.

I try adjusting the mirror, but it's not the alignment problem because I couldn't increase the efficiency.

So after the noise budget party with Jan and Frank, I check if the EOM works looking at the error signal from mixer out.  The laser frequency is scan at +/- 10V @ 100 Hz.

I'm not sure what is the corresponding freq span, but it must be less than 43 MHz(21.5 MHz x 2) because

three peaks of the signal could not be seen with +/- 10V span, so I use a voltage calibrator to slowly adjust the temperature and see the rest of the signals.

Nevertheless, the signals are there, so I try to lock the PMC again, and now the efficiency back to almost 80% again ( I have to re align again because of the earlier adjustment. The mixer out channel is monitored when I set the PMC gain to make sure there will be no oscillation.

I'm not sure what happen, loosen connectors, mode hopping in the laser, etc. I'll see if I can track this down , otherwise we could not have a stable locking system.

ELOG V3.1.3-